OpenStudy (anonymous):

What is the slope for y=-x? I don't know if its 0 or -1... Help! Please

sam (.sam.):

y=-x y=mx+c m is the slope

OpenStudy (anonymous):

In this case you have y=(-1)x so your m value is -1.

OpenStudy (anonymous):

y=mx+b y=-1x+0 (m is the slope) m=-1

OpenStudy (anonymous):

Oh okay! Thank you! Bc I knew the intercept was 0 but I forgot how to plug it in! Thanks again guys!:)

OpenStudy (anonymous):

no prob:)